Circular No. 7002

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 T. L. Cline, Goddard Space Flight Center (GSFC); and E. P. Mazets

and S. V. Golenetskii, Ioffe Physical-Technical Institute, St. Petersburg, write: “A > 50-keV transient of intensity comparable to a strong γ-ray burst triggered the GRB monitor on the Global Geo-Science Wind spacecraft on Aug. 27.432148 UT, exhibiting a persistent periodicity of 5.15 ± 0.02 s. This is the first γ-ray transient seen to clearly exhibit a periodic response in an uncollimated GRB detector since the 1979 March 5 event, itself the progenitor of the first known soft γ-ray repeater (SGR). This transient is assumed to originate from SGR 1900+14, in view of its current state of activity (see, e.g., IAUC 6929), although there is no directional information from the Konus instruments other than a north-ecliptic-hemisphere indication, consistent with SGR 1900+14. Since the GGS-Wind spacecraft is not in a near-earth orbit, comparison with other observations may be able to provide a source directional annulus and confirmation of this identity.”

R. Remillard, D. A. Smith, and A. Levine, Massachusetts Institute of Technology (MIT), report on behalf of the RXTE/ASM teams at MIT and at NASA/GSFC: “The RXTE All Sky Monitor has detected an x-ray flare at a location consistent (± 0^{o}.2) with the position of the candidate x-ray counterpart for SGR 1900+14 (Hurley et al. 1996, Ap.J. 463, L13). Two consecutive detections were made in each of two cameras on Aug. 27, showing a 2–12-keV intensity of 78 ± 12 mCrab at Aug. 27.5174 UT and 114 ± 11 mCrab at Aug. 27.5185. During the preceeding week, there were no significant detections made, and on Aug. 26, the average of 13 measurements implies a 3σ upper limit of 14 mCrab. On the day following the flare, seven measurements indicate a mean intensity 18 ± 5 mCrab. During all of the 90-s exposures associated with ASM detections, the timeseries data indicate steady emission with no evidence of SGR bursts.”

V592 HERCULIS E. O. Waagen, AAVSO, reports that this ‘tremendous outburst ampli-

tude dwarf nova’ (cf. Howell et al. 1995, Ap.J. 439, 337) is in outburst for only the third time since 1968, as indicated by the following visual magnitude estimates: Aug 24.901 UT, [15.3 (G. Poyner, Birmingham, England); 25.90, [13.2 (P. Schmeer, Bischmisheim, Germany); 26.835, 12.0 (T. Kinnunen, Espoo, Finland); 27.181, 13.0 (L. Shaw, Pinole, CA); 27.192, 13.0 (G. Hanson, Cave Creek, AZ); 28.051, 13.5 (J. Bortle, Stormville, NY); 28.233, 12.9 (Shaw).

1998 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green